### v1

@@ 4 Theoretical Insight
In this section, we present the theoretical insights underlying the proposed DF-GMVI method, including its covariance positivity-preserving and affine invariance properties. These insights are also applied to design adaptive time-stepping schemes and select gradient flows, which significantly enhance the performance of other Gaussian mixture variational inference methods (see Section 5.2).
A common issue with Gaussian mixture-based sampling methods is instability due to the covariance matrix becoming singular or non-positive. The following proposition establishes a condition on the time step $\Delta t$ that ensures the covariance matrix remains positive definite during the update process, which relies on our design of the quadrature rules used for approximating Hessian expectations in Section 3. The proposition guarantees that we can choose a sufficiently large $\Delta t$, enabling faster convergence without introducing numerical instability (see Section 5).
@@
(31) $\begin{split}C_{k}^{-1}(t+\Delta t)=&C_{k}^{-1}(t)+\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}}\\ =&C_{k}^{-1}(t)+\Delta t~{}\Bigl{(}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}-C_{k}^{-1}\\ &+\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}\Bigr{)}\\ =&(1-\Delta t)C_{k}^{-1}(t)+\Delta t~{}\Bigl{(}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}\\ &+\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}\Bigr{)}.\end{split}$
Since the matrix within the brackets is always positive semidefinite and $(1-\Delta t)C_{k}^{-1}(t)$ is positive definite, we conclude that $C_{k}^{-1}(t+\Delta t)$ is also positive definite.
Moreover, sampling methods are said to be affine invariant if their behavior remains unchanged under any invertible affine transformation. This property makes these methods especially effective for highly anisotropic posterior distributions [27, 21]. Their effectiveness arises from consistent performance across coordinate systems related by affine transformations. Specifically, the convergence properties of these methods can be understood by examining the optimal coordinate system, which minimizes anisotropy to the fullest extent, across all affine transformations. The following proposition establishes a restricted affine invariance of DF-GMVI method.
@@
The sampling algorithm applied to the posterior distribution, which is proportional to $\mathrm{e}^{-\Phi_{R}(\theta)}$, produces Gaussian mixture approximations parameterized by $\{m_{k}(t),C_{k}(t),w_{k}(t)\}_{k=1}^{K}$. When the algorithm is applied to the transformed posterior, with $\widetilde{\Phi}_{R}(\widetilde{\theta})=\Phi_{R}(T^{-1}(\widetilde{\theta}-d))$, under any invertible affine mapping $\varphi:\theta\rightarrow\widetilde{\theta}=T\theta+d$, where $T\in\mathcal{T}$ is an invertible matrix and $d$ is a translation vector, it produces updated parameters $\{\widetilde{m}_{k}(t),\widetilde{C}_{k}(t),\widetilde{w}_{k}(t)\}_{k=1}^{K}$. The algorithm is said to be $\mathcal{T}$-invariant if the following holds:
@@
We provide a proof of this proposition in Appendix B. The affine invariance is restricted because the Cholesky decomposition of the covariance is only invariant for lower triangular matrix group. The affine invariance enables the DF-GMVI method and variational inference methods based on natural gradient flow (20) to converge faster than other gradient flows with Gaussian mixture approximations, as further demonstrated numerically in Section 5.2.
@@ 5 Numerical Study
In this section, we present numerical studies regarding the proposed DF-GMVI method. We focus on posterior distributions of unknown parameters or fields arising in inverse problems that may exhibit multiple modes. Three types of model problems are considered:
@@
    A one-dimensional bimodal problem from [14]: we use this problem as a proof-of-concept example. Our result shows that the convergence rate remains unchanged no matter how overlapped the two modes are. Additionally, DF-GMVI outperforms the derivative free Gaussian mixture Kalman inversion presented in [14].
@@
    Several multi-dimensional problems: we use these problems to demonstrate that DF-GMVI is robust with respect to the posterior featuring multiple modes, infinite many modes, and modes with narrow and curved shapes. And the convergence rate remains robust across different problem dimensionalities. Additionally, we compare DF-GMVI with other state-of-the-art sampling approaches, including Gaussian mixture variational inference methods leveraging gradient or Hessian information, BBVI, and MCMC methods, to highlight its strengths. Notably, the theoretical insights presented in Section 4 are general and are used to design adaptive time-stepping schemes that significantly enhance the robustness of these comparable Gaussian mixture variational inference methods.
@@
For all tests, we use $\Delta t=0.5$ and run 200 iterations, resulting in a total time of $T=100$. We set $\alpha=10^{-3}$ in the derivative free quadrature rule in Definition 3.1. Our code is available online: https://github.com/PKU-CMEGroup/InverseProblems.jl/tree/master/Derivative-Free-Variational-Inference.
@@ 5.1 One-Dimensional Problems
The prior is $\rho_{\rm prior}\sim\mathcal{N}(3,2^{2})$ and consider different noise levels:
@@ 5.2 Multi-Dimensional Problems
In this subsection, we first examine several 2D sampling problems, along with their 100-dimensional modified versions. We then conclude with a comprehensive comparison to other state-of-the-art sampling approaches. For the 2D sampling problems, we define $\Phi_{R}(\theta)=\frac{1}{2}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)$, where $\theta=[\theta_{(1)},\theta_{(2)}]^{T}\in\mathbb{R}^{2}$, and $\mathcal{F}(\theta)$ is defined as follows.
@@ 1. Case A
    $\mathcal{F}(\theta)=y-A\theta\quad A=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\quad\textrm{ and }\quad y=\begin{bmatrix}0\\ 1\end{bmatrix}.$
@@ 2. Case B
    : The distribution has four modes with different weights [14, Appendix D]
@@ 3. Case C
    : The distribution is circular in shape [14, Appendix E], with its mass concentrated along an elliptical region,
    $\mathcal{F}(\theta)=2(y-\theta^{T}A\theta)\quad\textrm{ and }\quad A=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\quad y=\begin{bmatrix}1\end{bmatrix}.$
@@ 5. Case E
We apply DF-GMVI with $K=10,20$, and $40$ modes, each randomly initialized as $\mathcal{N}(0,I)$ with equal weights. The density estimations at the 200th iteration are shown in Fig. 2, with each row corresponding to a case (A to D). In each row, the reference density is displayed in the first panel, followed by the results from DF-GMVI with $K=10,20$, and $40$ modes. The estimated densities closely match the reference. The fourth panel in each row shows the error in terms of total variation over the iterations. DF-GMVI effectively captures multiple modes (Cases B and E), densities characterized by narrow, curved valleys (Cases D and E), which typically pose challenges for sampling methods, as well as densities with infinitely many maximum a posteriori points (Case C). The fourth panel demonstrates that DF-GMVI converges efficiently across all cases. However, it is important to note that when the posterior is Gaussian (Case A), DF-GMVI does not achieve exponential convergence due to the interactions among different Gaussian components; in contrast, Gaussian variational inference Eq. 13[13] achieves exponential convergence in this setting. Furthermore, when the number of modes $K$ is small, or when the initial Gaussian components are distant from the target mode, DF-GMVI may fail to capture all modes of the posterior density. Increasing $K$ helps mitigate this issue and enhances performance.
Figure 2: Multi-dimensional problems with dimensionality $2$, arranged from Case A to Case E from the top row to the bottom row. Each panel shows the reference density alongside the densities estimated by DF-GMVI with $K=10,20$ and $40$ (from left to right). The mean $m_{k}$ of each Gaussian component is marked by red circles, and the initial means are marked by gray crosses. The fourth panel shows the total variation distance between the reference density and the densities estimated by the DF-GMVI over the iterations.
@@
where $\theta\in\mathbb{R}^{2}$, $\theta^{c}\in\mathbb{R}^{N_{\theta}-2}$, and $K\in\mathbb{R}^{(N_{\theta}-2)\times 2}$ is an all-ones matrix. The function $\mathcal{F}$ is defined as in the previous 2-dimensional densities. These high-dimensional densities are constructed so that the marginal densities of $\theta$ remain computable and identical to that of the previous 2-dimensional densities, which serves as the error indicator for evaluating DF-GDVI.
Similarly, we apply DF-GMVI with $K=10,20$, and $40$ modes, each randomly initialized as $\mathcal{N}(0,I)$ with equal weights. The marginal densities obtained from the estimated 100-dimensional densities at the 200th iteration are shown in Fig. 3, using the same configuration as in Fig. 2. The estimated marginal densities closely align with the reference, and the total variation distances are similar to those from the 2-dimensional problems in Fig. 2. This confirms that the performance of DF-GMVI remains robust across different dimensionalities. We conjecture that the difficulty of the sampling problem depends primarily on the complexity of the target density (e.g., the number of modes) rather than the dimensionality.
Figure 3: Multi-dimensional problems with dimensionality 100, arranged from Case A to Case E from the top row to the bottom row. Each panel shows the marginal density of the reference alongside the marginal densities estimated by DF-GMVI with $K=10,20$, and $40$ (from left to right). The projected means of each Gaussian component are marked by red circles, and the projected initial means are marked by grey crosses. The fourth panel displays the total variation between the reference marginal density and the estimated marginal densities across the iterations.
Finally, we compare DF-GMVI with other state-of-the-art methods, including the Gaussian mixture approximation of the natural gradient flow (NGF-VI) Eq. 20 [41], its diagonal covariance variant (NGF-VI-D) [49], the Wasserstein gradient flow (WGF-VI)[40], BBVI [52], and the affine invariant MCMC method [27, 21]. NGF-VI, NGF-VI-D, and WGF-VI are not derivative free, requiring the computation of the gradient or Hessian of $\Phi_{R}$. Expectations are approximated using the mean-point approximation discussed in Appendix A. Motivated by the covariance positivity insight outlined in Section 4, for the NGF-VI and NGF-VI-D, we implement an adaptive time-stepping scheme defined as
@@
with $\Delta t_{\max}=0.5$ and $\beta=0.99$, ensuring that the covariance update $C_{k}^{-1}(t+\Delta t)=C_{k}^{-1}(t)+\Delta t\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}]$ maintains positive definiteness. For the WGF-VI, the covariance update equation can be written as
@@
We apply the Monte Carlo method with $5$ ensembles (same as DF-GMVI) to compute these integrals. We use empirical covariance instead of $C_{k}$ in the right-hand-side of the covariance update equation to reduce variance, and implement a similar adaptive time-stepping scheme defined as
@@
Fig. 4 shows the results obtained using Gaussian mixture variational inference approaches with $K=40$ and the affine invariant MCMC method with $J=10^{3}$ particles for the 2-dimensional problems. We observe that these variational inference approaches require very small time steps, particularly in Case E, to ensure stability. Our proposed adaptive time steppers significantly improve this. The BBVI method exhibits random noise, which can be mitigated by increasing the ensemble size for Monte Carlo integrals. Among the variational inference methods, those based on natural gradient flows with affine invariance outperform their counterparts based on Wasserstein gradient flows. Overall, the results obtained by DF-GMVI, as shown in Fig. 2, are similarly comparable and avoid the use of derivatives. In high dimensional spaces, the volume grows exponentially, which makes it difficult for random derivative free methods to explore the space efficiently. As a result, the performance of BBVI and derivative free MCMC methods typically deteriorates with increasing dimensionality, as illustrated in Fig. 5. However, the quadrature rule Definition 3.1 leverages gradient structure, allowing our DF-GMVI method to avoid this issue even in the 100-dimensional problem. Consequently, our DF-GMVI method outperforms the other approaches on the test problems presented in this subsection.
Figure 4: Comparison study for multi-dimensional problems with dimensionality $2$, arranged from Case A to Case E (from top to bottom). Each panel displays the reference density alongside the densities estimated by variational inference methods and the affine invariant MCMC (left to right: NGF-VI, NGF-VI-D, WGF-VI, BBVI, MCMC). The last figure shows the total variation distance between the reference density and the densities estimated by different methods over the iterations.
Figure 5: Comparison study for multi-dimensional problems with dimensionality $100$ (Case B). The figures display the marginal reference density and the estimated marginal densities obtained by BBVI with $K=40$ and ensemble sizes $J=64,256$, and $1024$ for Monte Carlo integration, and by affine invariant MCMC method with ensemble sizes $J=2^{10},2^{12}$, and $2^{14}$ (from left to right). The last figure shows the total variation distance between the marginal reference density and the estimated marginal densities over iterations.

### v3

@@ 5 Theoretical Insight
In this section, we present the theoretical insights underlying the proposed DF-GMVI method, including its covariance positivity preserving and affine invariance properties. These insights are also applied to the design of adaptive time-stepping schemes and select gradient flows, which significantly improve the performance of other Gaussian mixture variational inference methods (see Section 6.2).
A common issue with Gaussian mixture-based sampling methods is instability when using large time steps, due to the covariance matrix losing positive-definiteness, as discussed in Section 3. The following proposition establishes a condition on the time step $\Delta t$ that ensures the covariance matrix remains positive definite during the update process, which relies on our design of the quadrature rules used for approximating Hessian expectations in Section 4. The proposition guarantees that we can choose a sufficiently large $\Delta t$ (on the order of $\mathcal{O}(1)$), enabling faster convergence without introducing numerical instability (see Section 6).
@@
(18) $\begin{split}C_{k}^{-1}(t+\Delta t)=&C_{k}^{-1}(t)+\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}}\\ =&C_{k}^{-1}(t)+\Delta t~{}\Bigl{(}\sqrt{C_{k}(t)}^{-T}(6\mathrm{Diag}(A_{k}^{T}A_{k})+B_{k}^{T}B_{k})\sqrt{C_{k}(t)}^{-1}-C_{k}^{-1}(t)\\ &+\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}\Bigr{)}\\ =&(1-\Delta t)C_{k}^{-1}(t)+\Delta t~{}\Bigl{(}\sqrt{C_{k}(t)}^{-T}(6\mathrm{Diag}(A_{k}^{T}A_{k})+B_{k}^{T}B_{k})\sqrt{C_{k}(t)}^{-1}\\ &+\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}\Bigr{)}.\end{split}$
Since the matrix within the brackets is always positive semi-definite and $(1-\Delta t)C_{k}^{-1}(t)$ is positive definite, we conclude that $C_{k}^{-1}(t+\Delta t)$ is also positive definite.
Moreover, the following proposition establishes the affine invariance of the natural gradient flow (6), and a restricted affine invariance of the DF-GMVI method.
@@
The sampling algorithm applied to the posterior distribution, which is proportional to $\mathrm{e}^{-\Phi_{R}(\theta)}$, generates Gaussian mixture approximations parameterized by $\{m_{k}(t),C_{k}(t),w_{k}(t)\}_{k=1}^{K}$. For the transformed posterior, with $\widetilde{\Phi}_{R}(\widetilde{\theta})=\Phi_{R}(T^{-1}(\widetilde{\theta}-d))$, under an arbitrary invertible affine mapping $\varphi:\theta\rightarrow\widetilde{\theta}=T\theta+d$, where $T\in\mathcal{T}$ is an invertible matrix and $d$ is a translation vector, it produces updated parameters $\{\widetilde{m}_{k}(t),\widetilde{C}_{k}(t),\widetilde{w}_{k}(t)\}_{k=1}^{K}$. The algorithm is $\mathcal{T}$-invariant if the following holds:
@@
We provide a proof of this proposition in Appendix A. The affine invariance is restricted because the Cholesky decomposition of the covariance is only invariant for the lower triangular matrix group. The affine invariance enables the DF-GMVI method and variational inference methods based on natural gradient flow (6) to exhibit faster convergence than other gradient flows with Gaussian mixture approximations, as further demonstrated numerically in Section 6.2.
@@ 6 Numerical Study
In this section, we present numerical studies on the proposed DF-GMVI method. We focus on posterior distributions of unknown parameters or fields arising in inverse problems that may exhibit multiple modes. Three types of model problems are considered:
@@
    A one-dimensional bimodal problem from [15]: we use this problem as a proof-of-concept example. Our result shows that the convergence rate remains unchanged no matter how overlapped the two modes are. Moreover, DF-GMVI outperforms the derivative free Gaussian mixture Kalman inversion presented in [15].
@@
    Several multi-dimensional problems: we use these problems to demonstrate that DF-GMVI is robust with respect to the posterior featuring multiple modes, infinitely many modes, and modes with narrow and curved shapes. The convergence rate remains robust across different problem dimensionalities. Additionally, we compare DF-GMVI with other state-of-the-art sampling approaches, including Gaussian mixture variational inference methods leveraging gradient or Hessian information, BBVI, and MCMC methods, to highlight its strengths. Here for these comparable methods, we design adaptive time-stepping schemes, following our guideline in Section 3 to ensure covariance positivity. This strategy also substantially enhances the robustness of these Gaussian mixture variational inference methods.
@@
For all tests, we set $\Delta t=0.5$ and run 200 iterations, resulting in a total time of $T=100$. We set $\alpha=10^{-3}$ in the derivative free quadrature rule in Definition 4.1. A parameter sensitivity study is presented in Appendix B. Our code is available online: https://github.com/PKU-CMEGroup/InverseProblems.jl/tree/master/Derivative-Free-Variational-Inference.
@@ 6.1 One-Dimensional Problems
The prior is $\rho_{\rm prior}\sim\mathcal{N}(3,2^{2})$ and examine different noise levels:
@@ 6.2 Multi-Dimensional Problems
In this subsection, we first investigate several 2D sampling problems, along with their 100-dimensional modified versions. We then conclude with a comprehensive comparison to other state-of-the-art sampling approaches. For the 2D sampling problems, we define $\Phi_{R}(\theta)=\frac{1}{2}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)$, where $\theta=[\theta_{(1)},\theta_{(2)}]^{T}\in\mathbb{R}^{2}$, and $\mathcal{F}(\theta)$ is defined as follows.
@@ 1. Case A
    $\mathcal{F}(\theta)=y-A\theta\quad A=\begin{bmatrix}1&1\\ 1&2\end{bmatrix},\quad\textrm{ and }\quad y=\begin{bmatrix}0\\ 1\end{bmatrix}.$
@@ 2. Case B
    : The distribution has four modes with different weights [15, Appendix D].
@@ 3. Case C
    : The distribution is circular in shape [15, Appendix E] and contains infinitely many modes, as also used in Section 3,
    $\mathcal{F}(\theta)=\frac{y-(\theta_{(1)}^{2}+\theta_{(2)}^{2})}{0.3}\quad\textrm{ and }\quad\quad y=1.$
@@ 5. Case E
We apply DF-GMVI with $K=10,20$ and $40$ modes, each randomly initialized as $\mathcal{N}(0,I)$ with equal weights. The density estimations at the 200th iteration are shown in Fig. 3, with each row corresponding to a case (A to E). In each row, the reference density is displayed in the first panel, followed by the results from DF-GMVI with $K=10,20$ and $40$ modes. The estimated densities closely match the reference. The fifth panel in each row shows the error in terms of total variation over the iterations. DF-GMVI effectively captures multiple modes (Cases B and E), densities characterized by narrow, curved valleys (Cases D and E), which typically pose challenges for sampling methods, as well as densities with infinitely many maximum a posteriori points (Case C). The fifth panel demonstrates that DF-GMVI converges efficiently among all cases. However, it is important to note that when the posterior is Gaussian (Case A), DF-GMVI does not achieve exponential convergence due to the interactions among different Gaussian components; in contrast, Gaussian variational inference [14] achieves exponential convergence in this setting. Furthermore, when the number of modes $K$ is small, or when the initial Gaussian components are distant from the target mode, DF-GMVI may fail to capture all modes of the posterior density. Increasing $K$ helps mitigate this issue and enhances performance.
Figure 3: Multi-dimensional problems with dimensionality $2$, arranged from Case A to Case E from the top row to the bottom row. Each panel shows the reference density alongside the densities estimated by DF-GMVI with $K=10,20$ and $40$ (from left to right). The mean $m_{k}$ of each Gaussian component is marked by red circles, and the initial means are marked by gray crosses. The fifth panel shows the total variation distance between the reference density and the densities estimated by the DF-GMVI over the iterations.
@@
where $\theta\in\mathbb{R}^{2}$, $\theta^{c}\in\mathbb{R}^{N_{\theta}-2}$, and $K\in\mathbb{R}^{(N_{\theta}-2)\times 2}$ is an all-ones matrix. The function $\mathcal{F}$ is defined as in the previous 2-dimensional densities. These high-dimensional densities are constructed so that the marginal densities of $\theta$ preserve computability and match exactly that of the previous 2-dimensional densities, allowing for exact error evaluation through 2D marginal densities to evaluate DF-GMVI.
Similarly, we apply DF-GMVI with $K=10,20$ and $40$ modes, each randomly initialized as $\mathcal{N}(0,I)$ with equal weights. The marginal densities obtained from the estimated 100-dimensional densities at the 200th iteration are shown in Fig. 4, using the same configuration as in Fig. 3. The estimated marginal densities closely align with the reference, and the total variation distances are similar to those from the 2-dimensional problems in Fig. 3. This confirms that the performance of DF-GMVI remains robust across different dimensionalities. We conjecture that the difficulty of the sampling problem depends primarily on the complexity of the target density (e.g., the number of modes) rather than the dimensionality.
Figure 4: Multi-dimensional problems with dimensionality 100, arranged from Case A to Case E from the top row to the bottom row. Each panel shows the marginal density of the reference alongside the marginal densities estimated by DF-GMVI with $K=10,20$, and $40$ (from left to right). The projected means of each Gaussian component are marked by red circles, and the projected initial means are marked by gray crosses. The fifth panel displays the total variation between the reference marginal density and the estimated marginal densities across the iterations.
Finally, we compare DF-GMVI with other state-of-the-art methods, including the Gaussian mixture approximation of the natural gradient flow (NGF-VI) Eq. 6 [50], its diagonal covariance variant (NGF-VI-D) [60], the Wasserstein gradient flow (WGF-VI)[48], BBVI [63], and the affine invariant MCMC method [31, 23]. NGF-VI, NGF-VI-D, and WGF-VI are not derivative free, requiring the computation of the gradient or Hessian of $\Phi_{R}$. Expectations are approximated using the mean-point approximation Eq. 7. Motivated by the covariance positivity insight outlined in Section 5, for the NGF-VI and NGF-VI-D, we implement an adaptive time-stepping scheme defined as
@@
with $\Delta t_{\max}=0.5$ and $\beta=0.99$, ensuring that the covariance update $C_{k}^{-1}(t+\Delta t)=C_{k}^{-1}(t)+\Delta t\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}]$ maintains positive definiteness. The adaptive time-stepping for NGF-VI significantly improves efficiency compared to the fixed time step used in Section 3. For the WGF-VI, the covariance update equation can be written as
@@
We apply the Monte Carlo method with $5$ ensembles (same as DF-GMVI) to compute these integrals. We use empirical covariance instead of $C_{k}$ in the right hand side of the covariance update equation to reduce variance, and implement a similar adaptive time-stepping scheme defined as
@@
Figure 5 shows the results obtained from Gaussian mixture variational inference approaches with $K=40$ components and the affine-invariant MCMC method using $J=10^{3}$ particles, applied to 2-dimensional problems over 500 iterations. We observe that these variational inference approaches require very small time steps, particularly in Case E, to ensure stability. Our proposed adaptive time steppers significantly improve this. The BBVI method exhibits random noise, which can be mitigated by increasing the ensemble size for Monte Carlo integrals. Among the variational inference methods, those based on natural gradient flows with affine invariance outperform their counterparts based on Wasserstein gradient flows. Overall, the results obtained by DF-GMVI, as shown in Figs. 3 and 5, are similarly comparable and avoid the use of derivatives. In high dimensional spaces, the volume grows exponentially, which makes it difficult for random derivative free methods to explore the space efficiently. As a result, the performance of BBVI and derivative free MCMC methods typically deteriorates with increasing dimensionality, as illustrated in Fig. 6. However, the quadrature rule Definition 4.1 leverages gradient structure, allowing our DF-GMVI method to avoid this issue even in the 100-dimensional problem. Consequently, our DF-GMVI method outperforms the other approaches on the test problems presented in this subsection.
Figure 5: Comparison study for multi-dimensional problems with dimensionality $2$, arranged from Case A to Case E (from top to bottom). Each panel displays the reference density alongside the densities estimated by variational inference methods and the affine invariant MCMC (left to right: NGF-VI, NGF-VI-D, WGF-VI, BBVI, MCMC). The last figure shows the total variation distance between the reference density and the densities estimated by these methods and DF-GMVI over the iterations.
Figure 6: Comparison study for multi-dimensional problems with dimensionality $100$ (Case B). The figures display the marginal reference density and the estimated marginal densities obtained by BBVI with $K=40$ and ensemble sizes $J=64,256$, and $1024$ for Monte Carlo integration, and by affine invariant MCMC method with ensemble sizes $J=2^{10},2^{12}$, and $2^{14}$ (from left to right) at the $500$-th iteration. The last figure shows the total variation distance between the marginal reference density and the estimated marginal densities over iterations.